OpenStudy (lolipop_drum312):

The equation of a line is x + 3y = 14. What is the y-intercept of the line? choose the answer below (the same as abcd) 3/14 -14 14/3 14

OpenStudy (lolipop_drum312):

i think a idk

OpenStudy (jackthegreatest):

Ok set y = 0

OpenStudy (jackthegreatest):

Y intercept is what does x equal when y is 0

OpenStudy (lolipop_drum312):

then we add it to somethin or.... i have no clue

OpenStudy (jackthegreatest):

So: X + 3(0) = 14

OpenStudy (jackthegreatest):

X+0=14

OpenStudy (jackthegreatest):

X=14

OpenStudy (lolipop_drum312):

so its 14?

OpenStudy (lolipop_drum312):

D*

OpenStudy (jackthegreatest):

Yes, but does it make sense to you now?

OpenStudy (lolipop_drum312):

yes thank you i have 3 more btw

OpenStudy (jackthegreatest):

Post and tag me

OpenStudy (jackthegreatest):

I am away for 5 min

OpenStudy (lolipop_drum312):

What is the slope of the line passing through the points (2, −5) and (4, 1)?

OpenStudy (lolipop_drum312):

and ok

OpenStudy (lolipop_drum312):

3 ​−45−45 ​ 2 ​ 5454 ​

OpenStudy (lolipop_drum312):

same as abcd

OpenStudy (lolipop_drum312):

5/4* dfor D

OpenStudy (lolipop_drum312):

i think B

OpenStudy (lolipop_drum312):

4/5 for B btw

OpenStudy (lolipop_drum312):

but idk B sounds cool what do you think am i wrong

OpenStudy (narutofan2016):

it would be c because \[3y \div 3= 14-x \div3\] y= 14-x over 3. y would then equal 14/3

OpenStudy (lolipop_drum312):

so we um... divide or..... it could be C? or B? idk

OpenStudy (lolipop_drum312):

oh you must mean for the first one thanks